A rectangle is 7 cm wide and 10 cm long. Find the length of a diagonal of the rectangle.
Answer by oscargut(2103)

A rectangle is 7 cm wide and 10 cm long. Find the length of a diagonal of the rectangle.
Using Pythagoras theorem a^2=b^2+c^2 where a is the diagonal of a right triangle so d^2=10^2+7^2=149 then d=sqrt(149)
R: d=sqrt(149)